# Template AASArticle with Examples: v6.3.1111Released on March, 1st, 2021

### v1

@@ 3 Synthetic Observables
In this section we show the results of our various radiative transfer calculations. In Section 3.1 we compare the 1D and 2D LTE synthetic spectra to reveal the multidimensional differences between the two calculations. In Section 3.2, we show the differences in synthetic spectra between the 1D LTE and 1D non-LTE calculations to determine the impact of higher accuracy physics for this model. We describe how we construct the pseudo-2D non-LTE observables in Section 3.3, as well as show the light curves for our models.
@@
We show in Figure 2 the difference in synthetic spectra between our 1D wedge and 2D full ejecta calculations in LTE at a selection of lines of sight at 15 days and 30 days after the time of explosion, the former of which is around the time of maximum $B$-band magnitude (see appendix for all lines of sight). The 2D deviation from the 1D counterpart, illustrated in red, is most significant at early times in addition to being dependent on line of sight. For example, the equatorial line of sight shows modest differences with a slightly redder spectrum in the 2D calculation in the near UV and visual at 15 days post-explosion. This color disagreement persists at southern lines of sight, including at wedge 11 where it is most significant, but weakens and reverses at northern lines of sight. Wedge 11 also sees the most notable change in the trough shape of its Ca II H and K absorption feature ($\sim 3800$). Additionally, there is a modest excess of flux redward of $6000$ for the 2D spectra at the southernmost lines of sight, while the northernmost lines of sight see a fairly significant lack of flux in this region, including the three northernmost viewing angles which show a considerable departure in the bump at $\sim 8500$. The average absolute difference in flux between the 1D and 2D spectra as a percentage of the total flux at 15 days is 10%, with a median of 8%.
At 30 days post-explosion, the differences in spectra between the 1D and 2D calculations are smaller in flux. The 1D and 2D spectra show remarkably few differences at most lines of sight north of the equatorial plane. A slight change in the redward edge of the feature around $5500$ can be seen growing in strength as the line of sight moves southward of the equatorial line of sight.
For a few southern wedges at late times, we find that the total outgoing luminosity from the 1D LTE calculation is higher than that from the 2D LTE at the corresponding viewing angle without adding any extra luminosity in the 1D calculation. This affects wedge 14 and 15 non-trivially as early as 25 days. This is primarily due to these wedges having amounts of radioactive material that are significantly higher than the average wedge, enabling outgoing luminosities that are higher than from the corresponding lines of sight of the 2D calculation where photons can be re-distributed latitudinally. We include wedge 14 in Figure 2 for completeness, but do not include them in following plots where late time observables are relevant. We detail how this affects our multi-dimensional correction to the non-LTE observables in more detail in Section 3.3. We leave strategies to unify the outgoing luminosity between the 1D and 2D LTE calculations for profiles that result in an excess of luminosity in 1D, such as scaling the energy deposition as we do in our non-LTE calculations (see Section 2.3), to future work. Additionally, we note that our boundary lines of sight, wedge 1 and 15 (see appendix), show somewhat disjoint results in comparison to the other lines of sight, which may be attributable to spurious symmetry effects in the hydronuclear and/or radiative transfer calculations. Excluding wedges 14 and 15, we find an average and median absolute difference in flux between the 1D and 2D spectra as a percentage of the total flux at 30 days of 9% and 7%, respectively.
@@
Figure 3: 1D spectra at a selection of lines of sight at 15 and 30 days after the explosion. Spectra from our 1D LTE calculations are shown in black while the green shading indicates how the 1D non-LTE spectra deviate from the 1D LTE spectra (i.e. the outer edge of the green shading is the 1D non-LTE spectra). The lines of sight are labeled in orange, corresponding with the wedges in Figure 1.
@@
In Figure 3, we again show 1D LTE spectra for several lines of sight at 15 and 30 days after explosion, but now with the differences with the 1D non-LTE spectra, illustrated in green. Importantly, the difference in 1D LTE and non-LTE spectra are, for the most part, greater than that between 1D and 2D LTE calculations. In other words, non-LTE effects are of greater consequence on the observables than those that arise from multidimensionality. At 15 days post-explosion, a significant discrepancy between the non-LTE and LTE spectra is in the UV. Northern wedges show higher flux blueward of the Ca II H and K absorption feature and particularly at wedge 2 where the emission is considerably larger for the non-LTE spectra. A notable aspect of this disagreement is the pair of adjacent peaks centered at $\sim 3000$ that is most easily observed in the non-LTE spectrum of wedge 2. This distinct feature also exists in the LTE spectra between wedges 5 and 8 and is less prominent at more southern lines of sight in both non-LTE and LTE. However, it is prominently produced in the non-LTE calculation for wedge 2 where it does not exist in the LTE spectra. Conversely, at equatorial and southern lines of sight, the non-LTE spectra show lower emission in the near-UV around $3500$ compared to the LTE spectra.
@@
Overall, the emission in the visual region of the spectra at 15 days post-explosion is lower in the non-LTE calculation, with the strength of the variation decreasing from the northernmost line of sight. The discrepancy between the two sets of spectra in this region fluctuates somewhat across wavelengths. For example, there appears to be fairly consistent agreement of the flux in the absorption feature at $\sim 5000$. Redward of this, the non-LTE spectra is similarly featured but suppressed, especially at northern and equatorial lines of sight. Blueward of the shared absorption feature at $\sim 5000$, the non-LTE spectra produce a prominent absorption feature at $4500$ that is not present in any of the LTE spectra. We attribute this to Si iii $\lambda 4560$, which appears in some observed SNe Ia (see Section 4.1). The non-LTE spectra also show higher UV emission, especially at northern and equatorial lines of sight, in addition to slightly higher peaks at $\sim 4000$. Additionally, the non-LTE spectra show a deeper Si ii $\lambda 6355$ absorption feature and less prominent bump in the near-IR at northern and equatorial lines of sight. The integrated absolute differences in flux between the 1D LTE and non-LTE spectra as a percentage of the total flux have an average and mean of 16% and 12%, respectively, at 15 days post-explosion.
@@
The difference in spectra between non-LTE and LTE calculations grows much greater as time increases beyond the maximum light, exemplified in the right panel of Figure 3 where the spectra are compared at 30 days post-explosion, and declines in strength from the northernmost wedge. The two sets of spectra are disjoint at most lines of sight across all wavelengths except around $5000$. The non-LTE spectra are significantly bluer than their LTE counterparts, with considerable excesses in UV and B flux and a flattened near-IR in all but the two southernmost lines of sight. There are also exclusive, prominent features in the non-LTE spectra at this time, including the pair of peaks centered at $\sim 3000$ (which are also observed at 15 days post-explosion in the non-LTE treatment) and troughs that are just blueward of the pre-existing peaks at 4000 and 4600 Å. Most lines of sight in non-LTE also exhibit sharper features between 5000 and 6400 Å. Aside from wedges 14 and 15, which suffer similar issues in the 1D LTE calculation as described in Section 3.1, the average and mean integrated absolute difference in flux between these two calculations as a percentage of the total flux are 76% and 82%, respectively, at 30 days post-explosion.
Figure 4: Temperature and ionization profiles for our three radiative transfer calculations for wedge 12 at 18 and 27 days past maximum light. The 2D LTE, 1D LTE, and 1D non-LTE calculations are shown with solid, dashed, and dotted lines, respectively. The temperature and Si, Fe, and Co mean ionization values are displayed in black, red, yellow, and green, respectively.
To elucidate the origin of the differences between the non-LTE and LTE spectra, we show in Figure 4 the temperature and ionization profiles for both of these treatments for wedge 12 near maximum light and nine days later. At low velocities beneath the photosphere at both times shown, the temperature and ionization states are very similar between the different calculations, indicating that LTE conditions are met within this region. At maximum light, the similarity in ejecta states for this wedge extends out to $\sim 14,000$ km s-1, outward of which the temperature and ionization diverges between the non-LTE and LTE calculations. At high velocities, the non-LTE calculation leads to much hotter and more ionized material, likely attributable to photoionization and other non-thermal processes. The state of the ejecta at a later time, shown in the bottom panel of Figure 4, is even more starkly different between non-LTE and LTE, with a notable increase in temperature and mean ionization state at high velocities in the non-LTE calculation.
@@
3.3 Construction of Pseudo-2D non-LTE Observables
As evidenced by Figures 2 and 3, both multi-dimensional and non-LTE aspects of the observables are consequential to varying degrees. As proper multidimensional, non-LTE calculations are generally unfeasible at this time, we concatenate our three sets of radiative transfer results to produce an approximated facsimile of such a calculation. We refer to this new set of synthetic observables as “pseudo-2D non-LTE”. Specifically, we take the differences in flux at each wavelength between the 1D and 2D LTE calculations (the red bands in Figure 2, i.e. the multidimensional correction) and apply them to the 1D non-LTE spectra at each time step for each line of sight. Likewise, for the theoretical photometry of our models, we take the flux difference between the 1D LTE and 2D LTE in each band and apply them to the 1D non-LTE to construct our pseudo-2D non-LTE light curves. In the following sections, we compare these pseudo-2D non-LTE synthetic observables to observations of SNe Ia spectra and light curves, in addition to a variety of population correlations.
Table 1: Wedge Details
LTE non-LTE
$10^{9}$ cm s-1 $10^{42}$ erg s-1 2D 1D 1D pseudo-2D
1 1.49 8.33 -18.94 -18.95 -18.77 -18.89
2 1.42 9.20 -19.06 -19.06 -18.81 -18.81
3 1.37 9.81 -19.13 -19.14 -18.87 -18.87
4 1.30 10.36 -19.18 -19.18 -18.91 -18.93
5 1.25 10.85 -19.22 -19.20 -18.94 -18.98
6 1.20 11.36 -19.25 -19.21 -18.98 -19.04
7 1.16 11.64 -19.27 -19.22 -19.01 -19.09
8 1.12 12.13 -19.29 -19.22 -19.03 -19.12
9 1.11 12.49 -19.31 -19.23 -19.06 -19.15
10 1.09 12.67 -19.32 -19.24 -19.08 -19.16
11 1.01 12.93 -19.33 -19.22 -19.09 -19.21
12 1.05 13.15 -19.33 -19.20 -19.09 -19.22
13 1.17 13.37 -19.33 -19.29 -19.15 -19.20
14 1.19 13.70 -19.33 -19.28 -19.15 -19.20
15 1.17 13.85 -19.32 -19.35 -19.17 -19.13
@@ 3.3 Construction of Pseudo-2D non-LTE Observables
We show in Figure 5 the light curves for our four sets of synthetic observables for a representative selection of viewing angles. The bolometric light curves for a given line of sight are effectively identical due to our imposed constraint on the non-LTE and 1D LTE calculations that they match the bolometric output from the 2D calculations. However, some bolometric light curves appear slightly offset in Figure 5 due to small shifts in the time of maximum light in the $B$-band between the sets of observables. The change in time of $B$-band peak is no more than 1.6 d for all lines of sight outside of the northern boundary line of sight, with an average difference of 1.2 d between the 1D LTE and 1D non-LTE light curves, the largest shift between any corresponding set of models.
@@
The multidimensional effect in LTE on the light curves can be determined by comparing the yellow and red lines in Figure 5. This effect is relatively minor in bands $U$, $B$, $V$, and $R$, where the light curve shapes between the 1D and 2D LTE are very similar, with notable magnitude differences observed only in the $B$-band at later times of the southern line of sight. The 1D calculation does a much poorer job of reproducing the $I$-band, however, as the number and location of the local extrema vary non-trivially between 1D and 2D LTE.
@@
As indicated by the differences between the spectra in Figure 3, there is a dramatic change between the LTE and non-LTE light curves, especially relative to the aforementioned minor differences from dimensionality. While the $U$-band peaks at a somewhat similar magnitude between LTE and non-LTE treatments, the declines see significantly different shapes and rates, especially at northern latitudes (e.g. the top panel of Figure 5). The difference in peak magnitudes is more significant in the $B$-band and additionally varies slightly by line of sight. Crucially, the shape of the $B$-band light curve is starkly different in the non-LTE calculation with a consistently shallower decline from peak to 15 days post. We discuss this change in $B$-band light curve, including its relevance to the Phillips relation in particular, below in Section 4.2.
@@
In longer wavelength bands, we observe lower emission in the non-LTE light curves compared to LTE for most times between 0 and 15 days after maximum light. For the $V$- and $R$-band, this difference is heavily dependent on line of sight, as shown in Figure 5, where the peaks have a larger disparity at more northern lines of sight. The shapes of the light curves also vary between the radiative transfer calculations, especially in $V$-band where the decline is shallower in non-LTE. Additionally, we see a significant change in the post-peak $I$-band light curves, with the non-LTE exhibiting a mostly monotonic decline compared to the secondary local maxima seen in the LTE. This lack of secondary maximum in the non-LTE $I$-band light curves may be attributable to the stunted recombination of IGE elements in these calculations (see Figure 4) (Kasen, 2006). We note the relatively large discrepancy between the 1D and pseudo-2D non-LTE in the $I$-band, which corresponds to the disparity between the 1D and 2D LTE light curve in this band.

### v2

@@ 3 Synthetic Observables
Table 1: Wedge Details
$10^{9}$ cm s-1 $10^{42}$ erg s-1 mag
LTE non-LTE
2D 1D 1D pseudo-2D
1 1.49 8.33 -18.94 -18.86 -18.77 -18.87
2 1.42 9.20 -19.06 -19.06 -18.81 -18.81
3 1.37 9.81 -19.13 -19.13 -18.87 -18.88
4 1.30 10.36 -19.18 -19.16 -18.91 -18.95
5 1.25 10.85 -19.22 -19.18 -18.94 -19.00
6 1.20 11.36 -19.25 -19.19 -18.98 -19.05
7 1.16 11.64 -19.27 -19.19 -19.01 -19.10
8 1.12 12.13 -19.29 -19.19 -19.03 -19.14
9 1.11 12.49 -19.31 -19.20 -19.06 -19.17
10 1.09 12.67 -19.32 -19.22 -19.08 -19.18
11 1.01 12.93 -19.33 -19.20 -19.09 -19.22
12 1.05 13.15 -19.33 -19.19 -19.09 -19.23
13 1.17 13.37 -19.33 -19.28 -19.15 -19.20
14 1.19 13.70 -19.33 -19.28 -19.15 -19.20
15 1.17 13.85 -19.32 -19.35 -19.17 -19.14
In the follow subsections, we show the results of our various radiative transfer calculations. In Section 3.1 we compare the 1D and 2D LTE synthetic spectra to reveal the multidimensional differences between the two calculations. In Section 3.2, we show the differences in synthetic spectra between the 1D LTE and 1D non-LTE calculations to determine the impact of higher accuracy physics for this model. Finally, we compare the light curves for our sets of observables in 3.3.
@@
We show in Figure 2 the difference in synthetic spectra between our 1D wedge and 2D full ejecta calculations in LTE at a selection of lines of sight at 15 days and 30 days after the time of explosion, the former of which is around the time of maximum $B$-band magnitude (see Figure 15 in the Appendix for all lines of sight). The 2D deviation from the 1D counterpart, illustrated in red, is most significant at early times in addition to being dependent on line of sight. For example, the equatorial line of sight shows modest differences with a slightly redder spectrum in the 2D calculation in the near UV and visual at 15 days post-explosion. This color disagreement persists at southern lines of sight, including at wedge 11 where it is most significant, but weakens and reverses at northern lines of sight. Wedge 11 also sees the most notable change in the trough shape of its Ca ii H&K absorption feature ($\sim 3800$). Additionally, there is a modest excess of flux redward of $6000$ for the 2D spectra at the southernmost lines of sight, while the northernmost lines of sight see a fairly significant lack of flux in this region, including the three northernmost viewing angles which show a considerable departure in the Ca ii triplet at $8500$. The average absolute difference in flux between the 1D and 2D spectra as a percentage of the total flux at 15 days is 11%, with a median of 10%.
At 30 days post-explosion, the differences in spectra between the 1D and 2D calculations are smaller in flux. The 1D and 2D spectra show remarkably few differences at most lines of sight. For our boundary wedges (1 and 15), we find non-trivial differences between the 1D and 2D LTE spectra at later times, which may be attributable to spurious symmetry effects in the hydronuclear and/or radiative transfer calculations. Wedges 13 and 14 also show notable deviations. This is likely due to the three southernmost wedges containing significantly more 56Ni than the other wedges, as a result of the 56Ni distribution from this typical double detonation model being centered in the southern hemisphere. We find an average and median absolute difference in flux between the 1D and 2D spectra as a percentage of the total flux at 30 days of 14% and 9%, respectively.
@@
Figure 3: 1D spectra at a selection of lines of sight at 15 and 30 days after the explosion. Spectra from our 1D LTE calculations are shown in black while the green shading indicates how the 1D non-LTE spectra deviate from the 1D LTE spectra (i.e. the outer edge of the green shading is the 1D non-LTE spectra). The lines of sight are labeled in orange, corresponding with the wedges in Figure 1. Notable spectral features are labeled in red.
@@
In Figure 3, we again show 1D LTE spectra for several lines of sight at 15 and 30 days after explosion, but now with the differences with the 1D non-LTE spectra, illustrated in green. Importantly, the difference in 1D LTE and non-LTE spectra are, for the most part, greater than that between 1D and 2D LTE calculations. In other words, non-LTE effects are of greater consequence on the observables than those that arise from multidimensionality.
At 15 days post-explosion, a significant discrepancy between the non-LTE and LTE spectra is in the UV. Northern wedges show higher flux blueward of the Ca ii H&K absorption feature and particularly at wedge 2 where the emission is considerably larger for the non-LTE spectra. A notable aspect of this disagreement is the pair of adjacent peaks centered at $\sim 3000$ , which is largely formed due to IGE line blanketing (see J. M. DerKacy et al. 2020), most easily observed in the non-LTE spectrum of wedge 2. This distinct feature also exists in the LTE spectra between wedges 5 and 8 and is less prominent at more southern lines of sight in both non-LTE and LTE. However, it is prominently produced in the non-LTE calculation for wedge 2 where it does not exist in the LTE spectra. Conversely, at equatorial and southern lines of sight, the non-LTE spectra show lower emission in the near-UV around $3500$ compared to the LTE spectra.
@@
Overall, the emission in the visual region of the spectra at 15 days post-explosion is lower in the non-LTE calculation, with the strength of the variation decreasing from the northernmost line of sight. The discrepancy between the two sets of spectra in this region fluctuates somewhat across wavelengths. For example, there appears to be fairly consistent agreement of the flux in the Fe ii absorption complex at $\sim 5000$. Redward of this, the non-LTE spectra is similarly featured but suppressed, especially at northern and equatorial lines of sight. Blueward of the shared absorption feature at $\sim 5000$, the non-LTE spectra produce a prominent absorption feature at $\sim 4400$ that is not present in any of the LTE spectra. We attribute this to Si iii $\lambda 4560$, which appears in some observed SNe Ia (see Section 4.1). The non-LTE spectra also show higher UV emission, especially at northern and equatorial lines of sight, in addition to slightly higher peaks on either side of the Si ii $\lambda 4130$ absorption feature at $\sim 4000$. Additionally, the non-LTE spectra show a deeper Si ii $\lambda 6355$ absorption feature and an altered near-IR Ca ii triplet at northern and equatorial lines of sight. Lastly, we observe a very slight absorption feature at $\sim 7500$ at some lines of sight in the non-LTE spectra, which may be O i (X. Zhao et al., 2016). The integrated absolute differences in flux between the 1D LTE and non-LTE spectra as a percentage of the total flux have an average and mean of 16% and 11%, respectively, at 15 days post-explosion.
@@
The difference in spectra between non-LTE and LTE calculations grows much greater as time increases beyond the maximum light, exemplified in the right panel of Figure 3 where the spectra are compared at 30 days post-explosion, and declines in strength from the northernmost wedge. The two sets of spectra are disjoint at most lines of sight across all wavelengths except in the Fe ii-dominated region around $5000$. The non-LTE spectra are significantly bluer than their LTE counterparts, with considerable excesses in UV and B flux and a flattened near-IR in all but the two southernmost lines of sight. Similar stark differences between CMFGEN and Sedona spectra around this phase have previously been shown in S. Blondin et al. (2022).
@@
The average and mean integrated absolute difference in flux between these two calculations as a percentage of the total flux are 69% and 74%, respectively, at 30 days post-explosion.
Figure 4: Temperature and ionization profiles for our three radiative transfer calculations for wedges 4 and 12 at 18 and 27 days post-explosion. The 2D LTE, 1D LTE, and 1D non-LTE calculations are shown with solid, dashed, and dotted lines, respectively. The temperature and Si, Fe, and Co mean ionization values are displayed in black, red, yellow, and green, respectively.
To elucidate the origin of the differences between the non-LTE and LTE spectra, we show in Figure 4 the temperature and ionization profiles for both of these treatments for wedges 4 and 12 near maximum light and nine days later. At low velocities ($\lesssim 10,000$ km s-1) beneath the photosphere for both wedges at maximum light (left panels of Figure 4), the temperature and ionization states are similar between the different calculations, indicating that LTE conditions are generally met within this region. There is some disparity between the Si ionization states at very low velocities around maximum light. However, this region is predominately composed of IGE, and thus the Si ionization at these depths is inconsequential.
The maximum light similarities in ejecta states for these wedges extends out to $\sim 15,000$ km s-1, outward of which the temperature and ionization diverge between the non-LTE and LTE calculations. This point of divergence is associated with the density discontinuity of the double detonation model, arising from the interface between the core and shell detonations. At these higher velocities where the shell detonation ashes are found, the non-LTE effects become dominant due to the optically-thin material, leading to much higher ionization. Additionally, a spurious spike in mean ionization can be observed in this density discontinuity region in the 1D LTE ionization profile of wedge 12, which may arise due to the 1D ejecta profile improperly encapsulating the density discontinuity, LTE issues, or a combination of both.
The general agreement between ionization states at low velocities is weakened by 27 days (right panels of Figure 4), with the non-LTE profiles showing moderately higher ionization, along with temperature profiles that have a much shallower decline than the LTE counterparts. This is largely due to the significant reduction in density throughout the ejecta by this time. At high velocities, the state of the ejecta is even more starkly different between non-LTE and LTE at this later time, with a notable increase in temperature and mean ionization state at high velocities in the non-LTE calculation, especially for wedge 12.
S. Blondin et al. (2022) show a similar comparison for a 1D SN Ia toy model in their Figure 7, though we note an appreciable difference in input model as well as plotted times. S. Blondin et al. (2022) find temperature and Co ionization profiles that similarly differ between Sedona and CMFGEN calculations at low velocities around maximum light. However, the comparison from S. Blondin et al. (2022) shows CMFGEN temperature and ionization profiles that decline monotonically around maximum light, which is likely attributable to their smooth and monotonic toy model, as opposed to our non-LTE profiles which see an increase at high velocities. At later times, the difference in profiles of our work more closely resembles that of S. Blondin et al. (2022).
[…]
3.3 Light Curves
@@
We show in Figure 5 the light curves for our four sets of synthetic observables for a representative selection of viewing angles. The bolometric light curves for a given line of sight are effectively identical due to our imposed constraint on the non-LTE and 1D LTE calculations that they match the bolometric output from the 2D calculations. However, some bolometric light curves appear slightly offset in Figure 5 due to small shifts in the time of maximum light in the $B$-band between the sets of observables. The change in time of $B$-band peak is no more than 1.8 d for all lines of sight, with an average difference of 1.0 d between the 1D LTE and 1D non-LTE light curves, the largest shift between any corresponding set of models.
@@
The multidimensional effect in LTE on the light curves can be determined by comparing the yellow and red lines in Figure 5. This effect is relatively minor in bands $U$, $B$, $V$, and $R$, where the light curve shapes between the 1D and 2D LTE are very similar, with slight disagreement in peak magnitudes in $B$ and $V$ for some wedges. The 1D calculation does a much poorer job of reproducing the $I$-band, however, as the number and location of the local extrema vary non-trivially between 1D and 2D LTE.
@@
As indicated by the differences between the spectra in Figure 3, there is a dramatic change between the LTE and non-LTE light curves, especially relative to the aforementioned minor differences from dimensionality. While the $U$-band peaks at a similar magnitude between LTE and non-LTE treatments, the declines see significantly different shapes and rates, especially at northern latitudes (e.g. the top panel of Figure 5). The peak $B$-band magnitudes vary somewhat between the non-LTE and LTE, depending on the line of sight. Crucially, the shape of the $B$-band light curve is starkly different in the non-LTE calculation with a consistently shallower decline from peak to 15 days post. We discuss this change in $B$-band light curve, including its relevance to the Phillips relation in particular, below in Section 4.2.
@@
In longer wavelength bands, we observe lower emission in the non-LTE light curves compared to LTE for most times between 0 and 15 days after maximum light. For the $V$- and $R$-band, this difference is heavily dependent on line of sight, as shown in Figure 5, where the peaks have a larger disparity at northern lines of sight. The shapes of the light curves also vary between the radiative transfer calculations, especially in $V$-band where the decline is shallower in non-LTE. Additionally, we see a significant change in the post-peak $I$-band light curves, with the non-LTE exhibiting a mostly monotonic decline compared to the secondary local maxima seen in the LTE. This lack of secondary maximum in the non-LTE $I$-band light curves may be attributable to the stunted recombination of IGE elements in these calculations (see Figure 4) (D. Kasen, 2006). We note the relatively large discrepancy between the 1D and pseudo-2D non-LTE in the $I$-band, corresponding to the disparity between the 1D and 2D LTE light curve in this band, which indicates this band is somewhat questionable in its interpretation. Future work will strive to reduce uncertainties in the $I$-band to better understand how both multidimensionality and non-LTE affect local extrema.